# Permutation twisted cohomology, remixed

### v2

@@ 5. The (re-)twisted cohomology ring
In particular, the image of the comparison map $\operatorname{comp}_{G}$ is as a set, precisely the image of $\prod\hat{\psi}$. Moreover, the image of $\prod\hat{\psi}$ is dense in $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$.
@@
Every prime $\mathscr{P}(H,\mathfrak{p})\in\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))$ (in the notation of [BG25]) is explicitly the product of the prime $\mathfrak{p}$ of $\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H})$ and the full rings $\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{K})$ with $K$ different from $H$. The map $\prod\hat{\psi}$ sends the prime ideal $\mathscr{P}(H,\mathfrak{p})\in V_{{G}/\!\!/{H}}\subseteq\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))$ to the ideal $\{f\in\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))\mid\left(\prod\hat{\psi}\right)(f)\in\mathscr{P}(H,\mathfrak{p})\}$, and we have $\left(\prod\hat{\psi}\right)(f)\in\mathscr{P}(H,\mathfrak{p})$ if and only if $\hat{\Psi}^{H}(f)\in\mathfrak{p}$. Verification of commutativity now follows from the fact that given any $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$, $H\leq G$, and $\mathfrak{p}\in V_{{G}/\!\!/{H}}$, we have $\operatorname{cone}(f)\not\in\mathscr{P}(H,\mathfrak{p})$ if and only if $\hat{\Psi}^{H}(f)\in\mathfrak{p}$. Indeed, it is well-known that the comparison map on the tt-category $\operatorname{D}_{b}({G}/\!\!/{H})$ is a homeomorphism $\operatorname{Spc}(\operatorname{D}_{b}({G}/\!\!/{H}))\xrightarrow{\cong}\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H}))$, sending a prime $\mathscr{P}(H,\mathfrak{p})$ to the prime $\mathfrak{p}=\{\zeta\mid\operatorname{cone}(\zeta)\not\in\mathfrak{p}\}\subset\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H},k)$. Thus, $\zeta\in\mathfrak{p}$ if and only if $\operatorname{cone}(\zeta)\not\in\mathscr{P}(H,\mathfrak{p})$, and setting $\zeta=\hat{\Psi}^{H}(f)$ shows the result.
For the final statement, Proposition 5.6 implies the kernel of $d$ is nilpotent, so the result follows by a standard algebraic geometry result. ∎
@@
In particular, the image of the prime $\mathscr{P}(H,\mathfrak{p})$ under the comparison map $\operatorname{comp}_{G}$ is equivalently
$\left(\prod\hat{\psi}\right)(\mathscr{P}(H,\mathfrak{p}))=\{f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\mid\hat{\Psi}^{H}(f)\in\mathfrak{p}\}.$
5.1. Twisted cohomology under localization and tt-functors
Next, we’ll extend the results of [BG25, Section 14] regarding localization.
Definition 5.9.
Let $H$ be a subgroup of $G$. Let $S_{H}\subset\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ be the multiplicative subset generated by all $\iota^{H}_{C}$, where $C$ runs over the irreducible endotrivials $C\in\mathcal{B}(G)$. We define a $\mathbb{Z}$-graded ring
as the twist-zero part of the localization of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ with respect to $S_{H}$. Explicitly, the homogeneous elements of $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$ consist of fractions $\frac{f}{g}$ where $f,g\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ with the same $\mathcal{B}(G)$-twist $q$, and $g$ is a product of “$H$-local isomorphisms” $g^{\prime}\colon k\to C[s]$ with $C\in\mathcal{B}(G)$. Thus, $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$ is $\mathbb{Z}$-graded by the shift only. The homological degree of $\frac{f}{g}$ is the difference sts-t between the shifts of $f$ and $g$. In particular, $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(1)\cong\operatorname{H}^{\scriptscriptstyle\bullet}(G)$, following directly from Definition 5.5.
@@ 5.1. Twisted cohomology under localization and tt-functors
Recall for a morphism $f$ in $\mathscr{K}(G)$ we write
We have that $\operatorname{open}(f)$ is the preimage of $\operatorname{comp}_{G}$ of the principle open $Z(f)^{c}=\{\mathfrak{p}\mid f\not\in\mathfrak{p}\}$, and is the open locus of $\operatorname{Spc}(\mathscr{K}(G))$ where $f$ is invertible.
Construction 5.11.
We can perform the central localization (see [Bal10a]) of the whole category $\mathscr{K}(G)$
$\mathcal{L}_{G}(H):=\mathscr{K}(G)[S_{H}^{-1}].$
In fact, this localization has idempotent completion $\mathscr{K}(G)|_{U(H)}$ by the same argument as in [BG25, Construction 14.12] or [Bal10a, Theorem 3.6]. Explicitly, the category $\mathcal{L}_{G}(H)$ is the Verdier quotient of $\mathscr{K}(G)$ by the tt-ideal $\langle\{\operatorname{cone}(g)\mid g\in S_{H}\}\rangle$. It has the same objects as $\mathscr{K}(G)$ and morphisms $x\to y$ of the form $\frac{f}{g}$ where $g\colon k\to C$ belongs to $S_{H}$ for $C$ an effective endotrivial and $f\colon x\to C\otimes y$ is any morphism in $\mathscr{K}(G)$ with the same twist $C$ as the denominator $g$. Moreover, the $\mathbb{Z}$-graded endomorphism ring $\operatorname{End}^{\scriptscriptstyle\bullet}_{\mathcal{L}_{G}(H)}(k)$ of the unit in $\mathcal{L}_{G}(H)$ is the $\mathbb{Z}$-graded ring $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$.
Construction 5.12.
Twisted cohomology $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is graded over a monoid over the form $\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)}$. The ring homomorphisms induced by tt-functors or localization will be homogeneous with respect to a certain homomorphism $\gamma$ on the corresponding grading monoids.
Let $H\leq G$ be a subgroup and consider the central localization $(-)_{U(H)}\colon\mathscr{K}(G)\twoheadrightarrow\mathcal{L}_{G}(H)$. Here, the morphisms $\iota^{H}_{C}$ become isomorphisms, yielding a homomorphism on the grading
$\gamma=\gamma_{U(H)}\colon\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)}\to\mathbb{Z}$
defined by $\gamma(s,q)=s+h_{k(q)}(H)$ and we obtain a ring homomorphism
homogeneous with respect to the homomorphism $\gamma_{U(H)}$.
We may also consider the effects of restriction and modular fixed points on twisted cohomology. Note $\mathcal{B}(G)$ is more complicated than $\mathcal{N}\subseteq\mathcal{B}(G)$, where $\mathcal{N}$ consists of endotrivials inflated from normal subgroups of $G$ of index $p$, so in general, the shifts $\gamma$ may be more difficult to compute in the case of restriction.
Construction 5.13.
Let $H\trianglelefteq G$ be a normal subgroup. The tt-functor $\Psi^{H}\colon\mathscr{K}(G)\to\mathscr{K}(G/H)$ maps effective endotrivials with h-mark at $G$ equal to 0 to effective endotrivials at $G/H$ equal to 0. In fact, since deflation preserves irreducibility of a representation, $\Psi^{H}$ sends elements of $\mathcal{B}(G)$ to $\mathcal{B}(G/H)$.
This defines a homomorphism of graded monoids
given by $\gamma(s,q)=(s,\overline{q})$ where $\overline{q}$ is given by the surjection $\mathbb{N}^{\mathcal{B}(G)}\twoheadrightarrow\mathbb{N}^{\mathcal{B}(G/H)}$ along the inclusion $\mathcal{B}(G/H)\hookrightarrow\mathcal{B}(G)$ induced from inflation. Therefore, modular fixed points defines a ring homomorphism $\Psi^{H}:\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G/H)$ homogeneous with respect to $\gamma_{\Psi^{H}}$.
Given a group homomorphism $\alpha:G^{\prime}\to G$, restriction along $\alpha$ also defines a tt-functor $\alpha^{*}:\mathscr{K}(G)\to\mathscr{K}(G^{\prime})$. This again defines a corresponding ring homomorphism $\alpha^{*}:\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G^{\prime})$ homogeneous with respect to $\gamma_{\alpha^{*}}$.
Note in this case if $\alpha$ is not surjective, then $\gamma_{\alpha^{*}}$ may not be constant on the shift $s$ and in general $\overline{q}$ will be difficult to compute. Conversely, if $\alpha$ is surjective, then $\alpha^{*}$ is simply inflation, which may be easily computed, as it is section of modular fixed points. It follows that the homomorphism $\Psi^{H}$ on twisted cohomology is split surjective.
Remark 5.14.
In particular, setting the functor $F$ equal to the localization functor $(-)|_{U(H)}$ obtains the following commutative square.
$\operatorname{Spc}(\mathcal{L}_{G}(H))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\operatorname{comp}_{\mathcal{L}_{G}(H)}}$$\scriptstyle{\operatorname{comp}_{G}}$
Here, the left hand vertical map is the usual twist-free comparison map from [Bal10a] for the tt-category $\mathcal{L}_{G}(H)$ and the $\otimes$-invertible $k[1]$.
We combine the functors as before to strengthen [BG25, Proposition 14.21].
@@
Let $H\trianglelefteq G$ be a normal subgroup. Then we have a commutative square
$V_{G/H}=\operatorname{Spc}(\operatorname{D}_{b}(k[G/H]))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}(G/H))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\hat{\psi^{H}}}$$\scriptstyle{\operatorname{comp}_{\operatorname{D}_{b}(k[G/H])}}$$\scriptstyle{\operatorname{comp}_{G}}$
In particular, the diagonal is injective.
@@
The proof follows essentially the same as the proof of [BG25, Proposition 14.21], except without the added assumption that $G/H$ is elementary abelian. For this, we use Proposition 4.17 which asserts the quotient $\mathscr{K}(G/H)\to\operatorname{D}_{b}(k[G/H])$ is the central localization $(-)|_{U(1)}$ for any group $G/H$. ∎
6. Injectivity of the comparison map
The crux of injectivity of the comparison map is showing that, the diagonal of Proposition 5.15 is injective when $H$ is not necessarily normal in $G$ as well.
@@ 6. Injectivity of the comparison map
Suppose for contradiction that $H_{2}\not\leq_{G}H_{1}$. By Corollary 4.11, there exists an effective endotrivial $C\in\mathcal{B}(G)$ for which $\Psi^{H_{2}}(\iota_{C}^{H_{1}})$ is the zero map, and vice versa. Now, consider the map $\hat{\psi}^{H_{2}}\colon V_{{G}/\!\!/{H_{2}}}\hookrightarrow\operatorname{Spc}(\mathscr{K}(G))$. We have:
@@
Therefore, $V_{{G}/\!\!/{H_{2}}}\cap\operatorname{open}(\iota^{H_{1}}_{C})=\emptyset$ in $\operatorname{Spc}(\mathscr{K}(G))$. On the other hand, $V_{{G}/\!\!/{H_{1}}}\subseteq\operatorname{open}(\iota^{H_{1}}_{C})$, since $\iota^{H_{1}}_{C}$ is invertible in $\mathscr{K}(G)/\mathfrak{m}_{H_{1}}$. Therefore, $\mathscr{P}\in\operatorname{open}(\iota^{H_{1}}_{C})$ but $\mathscr{Q}\not\in\operatorname{open}(\iota^{H_{1}}_{C})$, a contradiction. Thus $H_{2}\leq_{G}H_{1}$, and by symmetry $H_{1}=_{G}H_{2}$. Set $H:=H_{1}.$
@@
and this map is injective from Proposition 5.15. Otherwise, we have the following commutative diagram.
$V_{N_{G}(H)/H}$$\operatorname{Spc}(\mathscr{K}(N_{G}(H))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(N_{G}(H)))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\operatorname{Res}^{*}}$$\scriptstyle{\operatorname{comp}_{N_{G}(H)}}$$\scriptstyle{\operatorname{comp}_{G}}$$\scriptstyle{\operatorname{Res}^{*}}$
@@
is injective. We establish this as a separate lemma.
@@
$V_{N_{G}(H)/H}$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{0}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{1}))$$\cdots$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{n}))$}
Here, the horizontal arrows are induced by restriction and the downwards arrows are $\hat{\psi}^{H}$. We will prove this statement inductively by assuming that the map $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))$ is injective and showing that the composition $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1}))$ is injective, which then of course implies the next map $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1}))$ is injective as well. Injectivity of the base case $H_{0}=N_{G}(H)$ is established by Proposition 5.15.
Let $\mathfrak{p},\mathfrak{q}\in\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))$. We claim that $\operatorname{Spec}(\operatorname{Res}^{H_{i+1}}_{H_{i}})(\mathfrak{p})=\operatorname{Spec}(\operatorname{Res}^{H_{i+1}}_{H_{i}})(\mathfrak{q})$ if and only if the following property $(\ast)$ is satisfied:
@@
The converse is straightforward. Suppose there exists an $f\colon k\to C[s]\in\mathfrak{p}$ for which $^{g}f\not\in\mathfrak{q}$ for all $g\in H_{i+1}$. Then, the trace product of $H_{i+1}$-conjugates
@@
is a morphism $f^{\prime}\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1})$. Indeed, the tensor product of chain complexes
@@
has $kH_{i+1}$-module structure (one may see for instance that $C^{\prime}$ is the restriction of the tensor induced complex $\operatorname{Ten}^{H_{i+1}}_{H_{i}}C$ via the Mackey formula - while tensor induction is not defined in general up to homotopy, in this case the restriction is well-behaved up to homotopy), and the image of $k$ in $C^{\prime}$ is $H_{i+1}$-stable. By primality of $\mathfrak{p}$ and $\mathfrak{q}$, we have $f^{\prime}\in\mathfrak{p}$ but $f^{\prime}\not\in\mathfrak{q}$, as desired.
Now, we show that if $\mathfrak{p},\mathfrak{q}\in\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))$ satisfy $(\ast)$ and belong to the image of $V_{N_{G}(H)/H}$, then $\mathfrak{p}=\mathfrak{q}$. The statement is trivial if $H_{i+1}/H_{i}$ acts on $H_{i}$ trivially, so assume the action is nontrivial. Recall from Proposition 5.7 that in this case, $\mathfrak{p}=\rho(H,\overline{\mathfrak{p}})$ with $\overline{\mathfrak{p}}\in V_{N_{G}(H)/H}$, and we have $\mathfrak{p}=\{f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i})\mid\hat{\Psi}^{H}(f)\in\overline{\mathfrak{p}}\}$. Let $f\in\mathfrak{p}$. By Lemma 4.10, there exists an effective endotrivial chain complex $C$ of $kH_{i}$-modules and morphism $\iota^{H}_{C}$ such that $\hat{\Psi}^{H}(\iota^{H}_{C})=\operatorname{id}_{k}$ and $\hat{\Psi}^{H}({}^{g}(\iota^{H}_{C}))=0$ for all $g\in H_{i+1}\setminus H_{i}$. Therefore, $\iota^{H}_{C}\not\in\mathfrak{p}$ (since $\overline{\mathfrak{p}}\neq V_{N_{G}(H)/H}$), but for all $g\in H_{i+1}\setminus H_{i}$, we have $^{g}(\iota^{H}_{C})\in\varphi(H,0)\subseteq\mathfrak{p}$. We construct a morphism $f^{\prime\prime}\in\mathfrak{p}$ as follows,

### v3

@@ 5.2. An identification for the image of the comparison map
In particular, the image of the comparison map $\operatorname{comp}_{G}$ is as a set, precisely the image of $\hat{\psi}^{h}$.
@@
Commutativity follows directly commutativity of the diagrams of 5.10 for all subgroups $H\leq G$. The latter statement follows since the top and left arrows are bijections of sets. ∎
@@
Unwinding things, the image of the prime $\mathscr{P}(H,\mathfrak{p})\in\operatorname{Spc}(\mathscr{K}(G))$, with $\mathfrak{p}$ regarded as an element of $\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H}))\cong V_{{G}/\!\!/{H}}$, under the comparison map $\operatorname{comp}_{G}$ is precisely
$\hat{\psi}^{h}(\mathscr{P}(H,\mathfrak{p}))=\langle f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\mid\hat{\Psi}^{H}(f)\in\mathfrak{p}\rangle.$
While the comparison map is in general not surjective, we deduce that it has dense image. The proof of this mirrors a standard commutative algebra fact, but with a slight modification due to the grading.
Corollary 5.13.
The image of the comparison map, $\operatorname{im}(\operatorname{comp}_{G})$, has dense image in $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$.
Proof.
Let $f\in\operatorname{H}^{s,q}(G)$ be a homogeneous element of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. By [BG25, Theorems 7.1, 7.2], if $\hat{\Psi}(f)=0$, then $f$ is $\otimes$-nilpotent. We claim this fact is sufficient to prove that $\operatorname{im}(\operatorname{comp}_{G})=\operatorname{im}(\hat{\psi}^{h})$ is dense in $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$. Indeed, this follows by a standard argument. Recall that $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ has an open basis given by supports $\operatorname{supp}(f)$ of homogeneous $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. Therefore, for all homogeneous $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$, if $\operatorname{supp}(f)\neq\emptyset$, then $f\not\in\mathfrak{p}$ for some homogeneous $\mathfrak{p}\in\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$. Hence $f$ is not $\otimes$-nilpotent, and it follows by our fact that neither is $\hat{\Psi}(f)$, so there exists a prime $\mathfrak{q}\in\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))$ with $\hat{\Psi}(f)\not\in\mathfrak{q}$. Therefore $f\not\in\hat{\psi}^{h}(\mathfrak{q})$, so $\hat{\psi}^{h}(\mathfrak{q})\in\operatorname{supp}(f)$. Thus $\operatorname{im}(\hat{\psi}^{h})$ intersects every nonempty basic open set $\operatorname{supp}(f)$, as desired. ∎
@@
[…]
6. Injectivity of the comparison map
The crux of injectivity of the comparison map is showing the diagonal of 5.10 is injective when a subgroup $H\leq G$ is not necessarily normal. To show this, we exploit the fact that $p$-groups are nilpotent and induct up a subnormal chain of subgroups of $G$. We first make a seemingly straightforward but subtle observation.
@@ 6. Injectivity of the comparison map
Let $H\leq G$, $C^{?}$ be a chain complex of free $k\Gamma_{G}$-modules, and set $C:=C^{1}$. Then $f\in\operatorname{Hom}_{\mathscr{K}(H)}(k,\operatorname{Res}^{G}_{H}C[s])$ is $G$-stable if and only if there exists a $\hat{f}\in\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])$ satisfying $\operatorname{Res}^{G}_{H}\hat{f}=f$.
@@
If one knows that $f$ is $G$-stable on the nose, the result is clear, but $f$ may only be $G$-stable only up to homotopy. By choosing a subnormal sequence $H:=H_{0}\triangleleft H_{1}\triangleleft\cdots\triangleleft H_{n}=G$ with $H_{i+1}:H_{i}]=p$, it suffices to assume $[G:H]=p$, as the result can be inductively applied to the subnormal sequence. We show $f$ is homotopic to a $G$-stable chain complex homomorphism.
We have a direct sum decomposition $\operatorname{Res}^{G}_{H}C=D\oplus K$, where $D$ has no contractible direct summands. Let $p_{D}$ and $p_{K}$ denote projection onto $D$ and $K$ respectively followed by inclusion into $C$. If $f$ is $G$-stable in $\mathscr{K}(G)$, then by 5.3, for all $g\in G$, $f$ and $^{g}f$ agree on $D$, that is, $p_{D}\circ{}^{g}f=p_{D}\circ f$ as chain complex homomorphisms for all $g\in G$, where $p_{D}$ denotes projection onto $D$ followed by inclusion into $\operatorname{Res}^{G}_{H}C$. Of course, $p_{D}+p_{K}=\operatorname{id}$.
[…]
is homotopic to $f$. We note the exclusion of the term $1H\in[G/H]$ is purely for making computation easier, since $p_{K}\circ p_{D}=0$. We claim $\hat{f}$ is also a $G$-stable chain complex homomorphism. Indeed, for all $g\in G$, we compute for $p_{K}$ and $p_{D}$:
$p_{D}\circ{}^{g}\hat{f}=p_{D}\circ{}^{g}(p_{D}\circ f)+p_{D}\circ{}^{g}p_{K}\circ\left(\sum_{g^{\prime}\in[G/H]-gH}{}^{g^{\prime}}(p_{D}\circ f)\right)=p_{D}\circ{}^{g}(p_{D}\circ f)=p_{D}\circ\hat{f}$,
and
[…]
$=p_{K}\circ\hat{f}$.
@@
Suppose for contradiction that $H_{2}\not\leq_{G}H_{1}$. By Theorem 4.12, there exists an effective endotrivial $C\in\mathcal{B}(G)$ for which $\Psi^{H_{2}}(\iota_{C}^{H_{1}})$ is the zero map, and vice versa. Now, consider the map $\hat{\psi}^{H_{2}}\colon V_{{G}/\!\!/{H_{2}}}\hookrightarrow\operatorname{Spc}(\mathscr{K}(G))$. We have:
@@
Therefore, $V_{{G}/\!\!/{H_{2}}}\cap\operatorname{open}(\iota^{H_{1}}_{C})=\emptyset$ in $\operatorname{Spc}(\mathscr{K}(G))$. On the other hand, $V_{{G}/\!\!/{H_{1}}}\subseteq\operatorname{open}(\iota^{H_{1}}_{C})$, since $\iota^{H_{1}}_{C}$ is invertible in $\mathscr{K}(G)/\mathscr{M}(H_{1})$. Therefore, $\mathscr{P}\in\operatorname{open}(\iota^{H_{1}}_{C})$ but $\mathscr{Q}\not\in\operatorname{open}(\iota^{H_{1}}_{C})$, a contradiction. Thus $H_{2}\leq_{G}H_{1}$, and by symmetry $H_{1}=_{G}H_{2}$. Set $H:=H_{1}.$
@@
and this map is injective from 5.10. Otherwise, we have the following commutative diagram.
$V_{N_{G}(H)/H}$$\operatorname{Spc}(\mathscr{K}(N_{G}(H))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(N_{G}(H)))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\rho^{G}_{N_{G}(H)}}$$\scriptstyle{\operatorname{comp}_{N_{G}(H)}}$$\scriptstyle{\operatorname{comp}_{G}}$$\scriptstyle{\rho^{G}_{N_{G}(H)}}$
@@
is injective. We establish this as a separate lemma below. ∎
@@
$V_{N_{G}(H)/H}$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{0}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{1}))$$\cdots$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{n}))$$\scriptstyle{\rho_{H_{0}}^{H_{1}}}$$\scriptstyle{\rho_{H_{1}}^{H_{2}}}$$\scriptstyle{\rho_{H_{n-1}}^{H_{n}}}$
Here we identify $V_{N_{G}(H)/H}=\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}(N_{G}(H)/H,k))$ via 5.11, the horizontal arrows are induced by restriction, and the downwards arrows are each the map $\hat{\psi}^{H}$. We will prove this statement inductively by assuming that the map $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))$ is injective and showing that the composition $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1}))$ is injective, which then of course implies the next map $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1}))$ is injective as well. Injectivity of the base case $H_{0}=N_{G}(H)$ is established by 5.10.
Let $\mathfrak{p},\mathfrak{q}\in\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))$. We claim that $\rho_{H_{i}}^{H_{i+1}}(\mathfrak{p})=\rho_{H_{i}}^{H_{i+1}}(\mathfrak{q})$ if and only if the following property $(\ast)$ is satisfied for the pair $(\mathfrak{p},\mathfrak{q})$:
@@
The converse is straightforward. We show the contrapositive: suppose there exists an $f\colon k\to C[s]\in\mathfrak{p}$ for which $^{g}f\not\in\mathfrak{q}$ for all $g\in H_{i+1}$. Then, the ‘norm’ product of $H_{i+1}$-conjugates
@@
may be lifted to a morphism $f^{\prime}\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1})$. Indeed, the tensor product of chain complexes
@@
is up to homotopy in the image of restriction from $kH_{i+1}$. This follows from the Mackey formula for Borel-Smith functions, as $\operatorname{CF}_{b}(-)$ is a rational $p$-biset functor (see [BY07] - in particular it is a Mackey functor) so we have,
$\operatorname{Res}^{H_{i+1}}_{H_{i}}\operatorname{Ind}^{H_{i+1}}_{H_{i}}h_{C}=\sum_{g\in H_{i+1}/H_{i}}{}^{g}h_{C},$
and we have an identification of rational $p$-biset functors $\operatorname{CF}_{b}(-)\cong\operatorname{Pic}(\mathscr{K}(-))$ from [MIL25b, Section 5]. Since the image of $f^{\prime}$ in $C^{\prime}$ is $H_{i+1}$-stable, the ‘norm’ product $f^{\prime}\colon k\mapsto C^{\prime}$ is in the image of restriction by 6.1. Now by primality of $\mathfrak{p}$ and $\mathfrak{q}$, we have $f^{\prime}\in\mathfrak{p}$ but $f^{\prime}\not\in\mathfrak{q}$, which is what we wanted.
Now, we show that if $\mathfrak{p},\mathfrak{q}\in\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))$ satisfy $(\ast)$ and belong to the image of $V_{N_{G}(H)/H}$, then $\mathfrak{p}=\mathfrak{q}$. The statement is straightforward if $H_{i+1}/H_{i}$ acts trivially on $H_{i}$, so let us consider the case where the action is nontrivial. From 5.11, we have $\mathfrak{p}=\mathscr{P}(H,\overline{\mathfrak{p}})$ with $\overline{\mathfrak{p}}\in V_{N_{G}(H)/H}$, and we have $\mathfrak{p}=\{f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i})\mid\hat{\Psi}^{H}(f)\in\overline{\mathfrak{p}}\}$. Let $f\in\mathfrak{p}$. By Lemma 4.11, there exists an effective endotrivial chain complex $C$ of $kH_{i}$-modules and morphism $\iota^{H}_{C}$ such that $\hat{\Psi}^{H}(\iota^{H}_{C})=\operatorname{id}_{k}$ and $\hat{\Psi}^{{}^{g}H}(\iota^{H}_{C})=0$ for all $g\in H_{i+1}\setminus H_{i}$, noting that $g\not\in N_{G}(H)$ so $^{g}H\neq H$. For instance, one may take the endotrivial $C$ arising from the regular representation $kH_{i}$. Therefore, $\iota^{H}_{C}\not\in\mathfrak{p}$, but for all $g\in H_{i+1}\setminus H_{i}$, we have $^{g}(\iota^{H}_{C})\in\mathscr{P}(H,0)\subseteq\mathfrak{p}$. We construct a morphism $f^{\prime\prime}\in\mathfrak{p}$ as follows,